A capacitor of capacitance C is charged by a battery of emf V & then disconnected . The work done by an external agent to slowly insert a di-electric of di -electric strength k and half the length of capaciitor plates is.........
 

Grade:12

1 Answers

Khimraj
5 years ago
then new capacitance will be C/2 + KC/2 = (1+K)C/2
and charge willremain constant
so work done will be change in energy
= Q2/(1+K)C – Q2/2C
= (Q2/2C)(1-K/1+K)
and Q = CV
SO
W = (1-K/1+K)(CV2/2)
Hope it clears.